OpenStudy (anonymous):

Please help for medal!! A large rectangle has a length of 3 inches and a width of 0.3 inches. A smaller rectangle has a length of x³ inches and a width of .25x inches. COMPLETE THE FOLLOWING OBJECTIVES A - D. a) What is the area of the large rectangle? b) What is the area of the small rectangle? c) Write a polynomial expression that represents how much greater the area of the larger rectangle is than the area of the smaller rectangle. d) Rewrite this expression as a difference of squares.

OpenStudy (anonymous):

a= .9 atm i cant solve it all gtg -_- studyhall is over

OpenStudy (anonymous):

The area of the larger rectangle is length multiplied by the width\[=3*0.3=0.9 Square inches\] The area of the smaller rectangle \[=x^3*0.25x^4 Squear inches \] The difference in the areas of the larger and smaller and rectangles \[.9-.25*x^4 or (\sqrt[.9])^2-(.5x^2)^2\]

OpenStudy (anonymous):

Hi @ksa0924 how about objective c?